OpenStudy (anonymous):

The sum of three consecutive numbers is 57. What is the largest of these numbers? 18 19 20 21

OpenStudy (anonymous):

X+x+1+x+2=57 3x+3=57 3x=54 x= 54/3 x= 18 x+1= 19 x+2= 20 Therefore 20 is the largest

OpenStudy (anonymous):

Thank you!! Can you help with another?